# Finite-time convergence to an $\epsilon$-efficient Nash equilibrium in potential games

### v3

@@ 3.2 Convergence time
[…]
$\beta$ $\geq\frac{1}{\max\{\epsilon/2,\Delta\}}\log\left((A^{N}-A^{N}(\epsilon/2))\left(\frac{4}{\epsilon A^{N}(\epsilon/2)}-\frac{1}{A^{N}(\epsilon/2)}\right)\right).$ (6)
@@
for
$t\geq\frac{25N^{2}A^{5}}{16\pi^{2}}e^{4\beta}\bigg{(}\log\log A^{N}+\log\beta+2\log\frac{4}{\epsilon}\bigg{)}.$
In other words, after running $t=\tilde{\mathcal{O}}(N^{2}A^{5}(\frac{A^{N}}{\epsilon})^{1/\max\{\epsilon,\Delta\}})$ rounds of log-linear learning with $\beta=\frac{1}{\max\{\epsilon,\Delta\}}\log\frac{A^{N}}{\epsilon}$ the expected value of the potential of the joint action at time $t$ is $\epsilon$-optimal.
Theorem 3.1 provides the first finite-time convergence rates to an $\epsilon$-efficient NE in general potential games. In order to be $\epsilon$-close to an efficient NE $\beta$ must scale as $\Omega(\frac{1}{\max\{\epsilon,\Delta\}}\log\frac{A^{N}}{\epsilon})$. Furthermore, since the convergence time scales as $e^{4\beta}$ it grows polynomially in $\mathcal{A}$ and $1/\epsilon$ and exponentially in $N$. The exponential dependence on $N$ is unavoidable without further assumptions on the game. However, by deriving problem-dependent bounds we are the first to avoid the exponential dependence in $1/\epsilon$, see Table 1. Note that the case $\Delta=0$ is trivial as all action profiles are efficient NEs. In Appendix E, we conduct numerical experiments to showcase the performance of log-linear learning with $\beta$ specified as in Equation (6). To summarize, our experiments demonstrate that log-linear learning with such a $\beta$ converges to an $\epsilon$-efficient NE. In particular, it can be seen that the suboptimality gap $\Delta$ is an important feature to quantify the convergence time and that the convergence time increases polynomially in $1/\epsilon$ for $\epsilon<\Delta$.
{collect*}
mytable
Table 1: Convergence time of log-linear learning to an $\epsilon$-efficient NE.
@@
Routing game with $K$ vertices Cost functions of $\tilde{\mathcal{O}}(e^{\frac{N}{\epsilon}})$
(Asadpour and Saberi, 2009) degree at most $p$
@@
Potential game with interchangeable $\lambda$-Lipschitz continuous $\tilde{\mathcal{O}}(N(\frac{A\lambda}{\epsilon})^{\frac{A}{\epsilon}})$
players (Shah and Shin, 2010) potential function
Potential game with interchangeable $A\geq 4$ $\tilde{\mathcal{O}}(N(\frac{N^{A}}{\epsilon})^{\frac{1}{\max\{\epsilon,\Delta\}}})$
players (Corollary B.7)
Potential games (Theorem 3.1) $A\geq 4$ $\tilde{\mathcal{O}}(N^{2}A^{5}(\frac{A^{N}}{\epsilon})^{\frac{1}{\max\{\epsilon,\Delta\}}})$
Proof 3.2.
outline Here, we provide an outline of the proof, the full proof is deferred to Appendix B.1. The argument begins with the following decomposition:
$\mathbb{E}_{a\sim\mu^{t}}[\Phi(a)]$ $\geq\underbrace{\mathbb{E}_{a\sim\mu}[\Phi(a)]}_{\text{First term}}-2\underbrace{\|\mu^{t}-\mu\|_{TV}}_{\text{Second term}}\underbrace{\max_{a\in\mathcal{A}^{N}}\Phi(a)}_{\leq 1}$
First term: To control the first term above, we rely on the novel lemma below which shows that an action profile sampled from $\mu$ is in expectation $\epsilon/2$-optimal if $\beta$ is sufficiently large.
@@
$\beta$ $\geq\frac{1}{\max\{\epsilon,\Delta\}}\log\left((A^{N}-A^{N}(\epsilon))\left(\frac{1}{\epsilon A^{N}(\epsilon)}-\frac{1}{A^{N}(\epsilon)}\right)\right),$
@@
The proof of this lemma is provided in Appendix B.2. We use it to control the first term as:
$\mathbb{E}_{a\sim\mu}[\Phi(a)]\geq\max_{a\in\mathcal{A}^{N}}\Phi(a)-\epsilon/2$
when $\beta$ is defined as in Equation (6).
Second term: We can control the mixing time of log-linear learning using the bound of Lemma 2.2:
$\|\mu^{t}-\mu\|_{TV}\leq\epsilon/4$
for $t\geq\frac{1}{\rho(PP^{*})}(\log\log\frac{1}{\mu_{\min}}+2\log\frac{4}{\epsilon})$. We then bound $\rho(PP^{*})$ using the novel lemma below.
@@
Consider a Markov chain over state space $\mathcal{A}^{N}$ with $A\geq 4$. Assume that there exists $p_{\min},p_{\max}\in(0,1]$, such that the corresponding transition matrix $P\in\mathbb{R}^{A^{N}\times A^{N}}$ satisfies: [^5] Such a $p_{\min}\in(0,1]$ exists if the probability of transitioning from any state to a neighboring state is positive.
@@
where $\mathcal{N}(a)=\{\tilde{a}\in\mathcal{A}^{N}\>|\>\exists i\in[N]:\tilde{a}_{-i}=a_{-i}\}$. Then the log-Sobolev constant $\rho(PP^{*})$ of matrix $PP^{*}$ is lower bounded by:

### v4

@@ 3.2 Convergence time in general potential games
Consider a potential game with a potential function $\Phi:\mathcal{A}^{N}\rightarrow[0,1]$ and with $A\geq 4$. [^2] We assume $A\geq 4$ to bound the log-Sobolev constant in Lemma 4.2. For $\epsilon\in(0,1)$ and initial distribution $\mu^{0}$, assume that players adhere to log-linear learning with:
$\beta\!$ $\geq\!\frac{1}{\Delta}\!\log\left(\!\!\left(A^{N}\!\!-\!A_{*}^{N}\right)\!\!\left(\!\frac{2}{\epsilon A_{*}^{N}}\!-\!\frac{1}{A_{*}^{N}}\right)\!\right)\!.\!$ (5)
@@
for $t\geq\frac{25N^{2}A^{5}}{16\pi^{2}}e^{4\beta}\bigg(\log\log A^{N}+\log\beta+2\log\frac{4}{\epsilon}\bigg)$.
In other words, after $t=\tilde{\mathcal{O}}(N^{2}A^{5}(\frac{A^{N}}{\epsilon})^{1/\Delta})$ rounds of log-linear learning with $\beta=\tilde{\Omega}\left(\frac{1}{\Delta}\log\frac{A^{N}}{\epsilon}\right)$ the expected potential function value of the joint action at time $t$ is $\epsilon$-optimal. We provide a proof of this Theorem in Section 4.
Here, we discuss the result. Theorem 3.1 provides the first finite-time convergence rate to an $\epsilon$-efficient Nash equilibrium in general potential games. For $\epsilon\in(0,1)$, the convergence time grows polynomially in $A$ and $1/\epsilon$ and exponentially in $N$. We note that the exponential dependence on $N$ is unavoidable since the problem of finding an efficient Nash equilibrium in a potential game is NP-hard. This was shown for the integral multicast game and the fair cost-sharing game, which are instances of a potential game [14, Theorem 5] and [9, Theorem 9], respectively. Thus, the exponential dependence on $N$ in our bounds reflects the intrinsic complexity of the problem, rather than a limitation of the chosen learning rule or the convergence analysis. However, to our knowledge, we are the first to avoid exponential dependence on $1/\epsilon$ (see Table 1), which we achieve by introducing the problem-dependent constant $\Delta$. Similar suboptimality-based notions have been widely used in the stochastic multi-armed bandit literature, both for regret analysis [3] and best-arm identification [20]. While direct computation of $\Delta$ is infeasible without access to the potential function, it can often be estimated using domain knowledge or sampling-based techniques.
Table 1: Convergence of log-linear learning to $\epsilon$-efficient Nash equilibrium.
@@
Routing game with Cost functions of $\tilde{\mathcal{O}}(e^{\frac{N}{\epsilon}})$
$K$ vertices [2] degree at most $p$
@@
Potential game with inter- $\lambda$-Lipschitz $\tilde{\mathcal{O}}(N(\frac{A\lambda}{\epsilon})^{\frac{A}{\epsilon}})$
changeable players [38] continuous potential
Corollary 3.3 $A\geq 4$ $\tilde{\mathcal{O}}(N(\frac{N^{A}}{\epsilon})^{\frac{1}{\Delta}})$
Theorem 3.1 $A\geq 4$ $\tilde{\mathcal{O}}(N^{2}A^{5}(\frac{A^{N}}{\epsilon})^{\frac{1}{\Delta}})$
3.3 Convergence time in symmetric potential games
In the following, we additionally assume that the potential game is symmetric, that is, players are interchangeable. Note that many real-world examples, such as instances of resource allocation and coverage games [27], are symmetric.
Definition 3.2.
A game is symmetric if for any permutation $\pi$ of $\{1,\ldots,N\}$ it holds that:
$U_{i}(a_{1},\ldots,a_{N})=U_{\pi(i)}(a_{\pi(1)},\ldots,a_{\pi(N)}).$
In other words, a player’s utility depends solely on the number of players selecting each action and not on their identity. Thus, in a symmetric potential game, if $A<N$, the potential function $\Phi$ can be redefined in terms of a lower-dimensional function $\Phi_{m}:\Psi^{\mathcal{A}}\rightarrow[0,1]$, where:
with cardinality $Y=|\Psi^{\mathcal{A}}|\leq(N+1)^{A-1}$. Note that the cardinality of $\Psi^{\mathcal{A}}$ with $\mathcal{O}(N^{A})$ is smaller than that of the original state space $\mathcal{A}^{N}$ with $A^{N}$. At the same time, for any $a\in\mathcal{A}^{N}$, it holds that $\Phi(a)=\Phi_{m}(x(a))$, where $x(a)=(x_{1}(a),\ldots,x_{A}(a))$ and $x_{j}(a)$ denotes the fraction of players that selected action $j\in\mathcal{A}$, i.e., $x_{j}(a)=1/N|\{i\in[N]\>|\>a_{i}=j\}|$.
In the following, we assume all players follow the modified log-linear learning dynamics from [38]. In modified log-linear learning, a variant of log-linear learning, every player $i$ has an independent exponential clock with rate $\alpha/z_{i}^{t}$. Player $i$’s exponential clock is an exponentially distributed random variable of mean $\alpha/z_{i}^{t}$, where $\alpha>0$ is a parameter and $z_{i}^{t}:=1/N|\{j\in[N]\>|\>a_{j}^{t}=a_{i}^{t}\}|$ counts the number of players playing the same action as player $i$. When player $i$’s clock rings, i.e., the sampled $\exp(\alpha/z_{i}^{t})$ waiting time elapses, that player immediately resamples her action according to $p_{i}^{t}$ defined in Equation (2).
Modified log-linear learning induces an aperiodic and irreducible Markov chain on the lower-dimensional state space $\Psi^{\mathcal{A}}$ with stationary distribution $\mu_{m}\in\Delta(\Psi^{\mathcal{A}})$ given by [38, Lemma 2]:
$\mu_{m}(x)=\frac{e^{\beta\Phi_{m}}(x)}{\sum_{\tilde{x}\in\Psi^{\mathcal{A}}}e^{\beta\Phi_{m}}(\tilde{x})},\quad\forall x\in\Psi^{\mathcal{A}}.$
Corollary 3.3.
Consider a symmetric potential game with potential function $\Phi_{m}:\Psi^{\mathcal{A}}\rightarrow[0,1]$. For $\epsilon\in(0,1)$ and initial distribution $\mu^{0}$, assume that players adhere to modified log-linear learning with:
$\beta\geq$ $\frac{1}{\Delta}\log\bigg((N+1)^{A-1}\bigg(\frac{1}{\epsilon Y_{*}}-\frac{1}{Y_{*}}\bigg)\bigg),$
where $Y_{*}\!=\lvert\{x^{*}\in\Psi^{\mathcal{A}}\>|\>x^{*}\in\operatorname*{arg\,max}_{x\in\Psi^{\mathcal{A}}}\!\!\Phi_{m}(x)\}\rvert$ denotes the cardinality of the set of potential maximizers. Then,
$\mathbb{E}_{x\sim\mu^{t}}[\Phi_{m}(x)]\geq\max_{x\in\Psi^{\mathcal{A}}}\Phi_{m}(x)-\epsilon,$
for $t\geq\frac{N}{\alpha c}e^{3\beta}\left(\log((A-1)\log(N+1))+\log\beta+2\log\frac{4}{\epsilon}\right)$, where $c>0$ is some constant.
We provide the proof of the above corollary in Appendix B.2. The result states that after $t=\tilde{\mathcal{O}}(N(N^{A}/\epsilon)^{\frac{1}{\Delta}})$ rounds the expected potential function value at time $t$ is $\epsilon$-optimal. The polynomial dependence on $N$ crucially relies on considering exponential clocks with dynamic means of the form $\alpha/z_{i}^{t}$ in modified log-linear learning. In contrast, in classical log-linear learning, which considers exponential clocks with mean $1$, the convergence time depends exponentially on $N$ [38, Example 2]. Furthermore, if $A=\mathcal{O}(\log N/\log\log N)$, then the convergence time to an $\epsilon$-efficient Nash equilibrium is $\tilde{\mathcal{O}}(N(N^{\log N}/\epsilon)^{\frac{1}{\Delta}})$ which is quasi-polynomial in the input size $\log(N^{A})=\tilde{\mathcal{O}}(N^{\log N})$ and polynomial in $1/\epsilon$. Note that the assumption that the cardinality of the action space is exponentially smaller than the number of players is realistic in applications like routing games, where the number of routes is considerably smaller than the number of agents.
For symmetric potential games with a $\lambda$-Lipschitz-continuous potential function, [38] prove a convergence time of $\tilde{\mathcal{O}}(N(\frac{A\lambda}{\epsilon})^{\frac{A}{\epsilon}})$. Our result does not rely on this Lipschitz continuity assumption and significantly improves the dependence on $\epsilon$ from exponential to polynomial.
4 Proof of Theorem 3.1
In this section, we prove our main result Theorem 3.1. As we analyze the convergence time of log-linear learning by studying the associated Markov chain, we review the basic concepts and properties of Markov chains in Appendix A.
The proof leverages the following decomposition based on the Cauchy-Schwarz inequality:
$\mathbb{E}_{a\sim\mu^{t}}[\Phi(a)]$ $\geq\underbrace{\mathbb{E}_{a\sim\mu}[\Phi(a)]}_{\text{First term}}-2\underbrace{\|\mu^{t}-\mu\|_{TV}}_{\text{Second term}}\underbrace{\max_{a\in\mathcal{A}^{N}}\Phi(a)}_{\leq 1}.$
To control the first term, we propose a novel lemma, Lemma 4.1, that provides a lower bound on $\mathbb{E}_{a\sim\mu}[\Phi(a)]$ if $\beta$ is sufficiently large. The second term is related to the mixing time of log-linear learning. Thus, to control the second term, we leverage mixing-time bounds based on the log-Sobolev constant, where the log-Sobolev constant is defined in Equation (24) in Appendix A. We rely on another novel lemma, Lemma 4.2, to bound the log-Sobolev constant of the Markov chain induced by log-linear learning. Before we provide a formal proof of Theorem 3.1, we state the two lemmas mentioned above.
@@ 3.3 Convergence time in symmetric potential games
$\beta$ $\geq\frac{1}{\Delta}\log\left(\left(A^{N}-A_{*}^{N}\right)\left(\frac{1}{\epsilon A_{*}^{N}}-\frac{1}{A_{*}^{N}}\right)\right),$
@@
The proof of this lemma is provided in Appendix B.1.
@@
Consider a Markov chain $X_{t}$ over state space $\mathcal{A}^{N}$ with $A\geq 4$. Assume that there exists $p_{\min},p_{\max}\in(0,1]$, such that the corresponding transition matrix $P$ satisfies:
@@
where $\mathcal{N}(a)=\{\tilde{a}\in\mathcal{A}^{N}\>|\>\exists i\in[N]:\tilde{a}_{-i}=a_{-i}\}$. Then the log-Sobolev constant $\rho(PP^{*})$ of $PP^{*}$ is lower bounded by: